OpenStudy (anonymous):

i need help can someone help me with eliminating a variable to set up a quadratic equation

OpenStudy (amistre64):

if we knew the equation, maybe

OpenStudy (anonymous):

L=25ft W=10ft A=250sq ft

OpenStudy (amistre64):

those arent variables; since each one has a specific value.

OpenStudy (anonymous):

it says that A=LW and to define the relationship and then elimate a variable

OpenStudy (amistre64):

the relationship is defined as; The Area of a rectangule varies directly to the product of the length and width

OpenStudy (anonymous):

so what formula would i use to set up the quadratic equation

OpenStudy (amistre64):

Without knowing precisely what it is they are talking about, I wouldnt be able to give you a better answer

OpenStudy (anonymous):

1.Different pets need to have different areas. For example, a horse needs more space than a dog. Determine an area that is appropriate for your pet. 2.Define the shape of the rectangular area by establishing a relationship between the length and width of the rectangle. For example, L = 2W + 5, or W = 3L – 4. Be sure to include the appropriate units (inches, feet, yards, miles, or meters). 3.Using the fact that A = LW, together with the relationship defined in step 2, eliminate one of the variables to set up a quadratic equation. 4.Solve the quadratic equation using any of the techniques learned in this unit. The solution(s) will be one of the dimensions; use step 2 to find the other.

OpenStudy (amistre64):

perhaps I would assume this for an answer: L = 25; Width = 10; L = W+15 250 = L*W 250 = W(W+15) 250 = W^2 +15W w^2 +15W -250 = 0

OpenStudy (anonymous):

my L=25 and W=10 if helps

OpenStudy (amistre64):

If i read you correctly, I posted an adequate solution. But lately, that a pretty big if :)

OpenStudy (anonymous):

which means

OpenStudy (anonymous):

ok thank you i think i got it now sorry i missed the reply you posted